To check whether a number is divisible by 19 one can repeatedly add twice the last digit
to the rest of the number. For example:
13547→ 1354 + 7 ∗ 2 = 1368 → 136 + 8 ∗ 2 = 152 → 15 + 4 = 19
thus 13547 is divisible by 19 (indeed, 13547=19*713). Give a proof of this divisibility rule.
1
Expert's answer
2019-10-28T10:02:00-0400

Suppose we have the integer number "10a+b," where "b" is the digit of the units (this writing form is unique if we say that "b" is a digit).

The rule says to consider "a+2b."

Suppose this is a multiple of "19," namely "a+2b=19k," where "k" is some integer number.

Then "a=19k-2b" and hence


"10a+b=10(19k-2b)+b=190k-19b=19(10k-1)"

is a multiple of "19."

The condition is necessary too. Suppose "10a+b=19n" for some integer "n." Then "b=19n-10a" and hence


"a+2b=a+2(19n-10a)=38n-19a=19(2n-a)"

is a multiple of "19."